Find the product of ( 3-i 4,
-4 3+i )
and ( 2 5-2i
-5 - 2i 2 )
1
Expert's answer
2020-06-03T16:11:06-0400

Doing dot product of "((3- \frac{i}{4}),(\frac{-4}{3} + i))" and "((\frac{2}{5} - 2i),(\frac{-5-2i}{2}))"


Product = "((3- \frac{i}{4}),(\frac{-4}{3} + i))" * "((\frac{2}{5} - 2i),(\frac{-5-2i}{2}))"


=( "(3- \frac{i}{4})" * "(\frac{2}{5} - 2i)" , "(\frac{-4}{3} + i)" * "(\frac{-5-2i}{2})" )


solving it we get,


= "((\frac{7}{10} - \frac{61}{10}i) , (\frac{7}{3} - \frac{7}{6}i))"


since "i^{2} = -1"